# Cold molecular gas in the hot nuclear wind of the Milky Way

### v1

@@
Using the Large Millimeter Telescope and the SEQUOIA 3 mm focal plane array, we have searched for molecular line emission from two atomic clouds associated with the Fermi Bubble of the Milky Way. Neither 12CO nor 13CO J=1-0 emission is detected from the H i cloud, MW-C20. 12CO J=1-0 emission is detected from MW-C21 that is distributed within 11 clumps with most of the CO luminosity coming from a single clump. However, we find no 13CO emission to a 3 brightness temperature limit of 0.3 K. Using this limit and RADEX non-local thermodynamic equilibrium (non-LTE) excitation models, we derive H2 column density upper limits of (0.4-3)1021 cm-2 for a set of physical conditions and a H2 to 12CO abundance ratio of 104. Model CO-to-H2 conversion factors are derived for each set of physical conditions. We find the maximum value is 1.61020 cm-2/(K km s. Increasing [H2/12CO] to 105 to account for photodissociation and cosmic ray ionization increases the column density and upper limits by a factor of 10. Applying these limits to the CO luminosities, the upper limit to the total molecular mass in MW-C21 is 1322 M⊙, corresponding to 27% of the neutral gas mass. For the three clumps that are fully resolved, lower limits to the virial ratios are 28832, 6828, and 15739, which suggest that these structures are bound by external pressure to remain dynamically stable over the entrainment time of 2106 years or are being disrupted by shear and expansion over the clump crossing times of 3-8105 years. The observations presented in this study add to the growing census of cold gas entrained within the Galactic Center wind.
@@ 4.2 Clump masses
Using the RADEX models, an upper limit to the conversion factor, , can be derived that is consistent with the non-detection of 13CO emission. For each set of parameters, the model 12CO brightness temperature, TB(12CO), is calculated at the H2 column density upper limit. The model 12CO integrated intensity, , is the product of this brightness temperature and the line width of 11 km s-1. The upper limit to is =. The values for TB(12CO) and for each model are listed in Table 2.
@@ 4.3 Virial ratio
The molecular clumps are embedded within a cloud of atomic gas with a mass of 428 M⊙ distributed within a radius of 40 pc (Di Teodoro et al., 2018). Assuming a spherical cloud, the mean volume density is 0.06 cm-3. An upper limit to the gas kinetic temperature can be made assuming the measured HI 21 cm velocity dispersion from Di Teodoro et al. (2018) is due to purely thermal motions, such that =3.5104 K, where is the H i atomic mass and is the Boltzmann constant. The upper limit to the cloud thermal pressure, nT, is 2.2103 K cm-3= 4.410-13 dynes cm-2, which is insufficient to bound the internal pressure of the molecular clumps. Cosmic-ray pressure and magnetic field surface tension are possible bounding agents but the strength of each is not well defined in the local environment of MW-C21.
@@ 5 Conclusions
We have searched for CO emission from 2 atomic gas clouds associated with the Fermi Bubble. 12CO J=1-0 emission is detected in the cloud MW-C21. This emission is distributed in 11 discrete clumps. 12CO is not detected in the cloud MW-C20 and 13CO J=1-0 emission is not detected in either cloud. Based on RADEX models to predict 13CO J=1-0 brightness temperatures over a range of physical conditions, the absence of detected 13CO emission constrains the H2 column density to be less than 31021 cm-2 and the CO to H2 conversion factor, , to be 1.61020 cm-2(K km s-1)-1, assuming [H2/12CO]=104. Clump mass upper limits range from 1 M⊙ to 95 M⊙ with a total molecular mass 132 M⊙, which comprises 27% of the neutral gas mass in the cloud. Mass surface density upper limits for the three clumps that are well resolved by the 12″ beam are 18, 42, and 9 M⊙pc-2. The derived virial ratio upper limits for the three resolved clumps are high (288, 68, 157), yet molecular gas is still present over the entrainment time scale, which suggests that these molecular features are bound by an external pressure component.

### v2

@@
Using the Large Millimeter Telescope and the SEQUOIA 3 mm focal plane array, we have searched for molecular line emission from two atomic clouds associated with the Fermi Bubble of the Milky Way. Neither 12CO nor 13CO J=1-0 emission is detected from the H i cloud, MW-C20. 12CO J=1-0 emission is detected from MW-C21 that is distributed within 11 clumps with most of the CO luminosity coming from a single clump. However, we find no 13CO emission to a 3 brightness temperature limit of 0.3 K. Using this limit and RADEX non-local thermodynamic equilibrium (non-LTE) excitation models, we derive H2 column density upper limits of (0.4-3)1021 cm-2 for a set of physical conditions and a H2 to 12CO abundance ratio of 104. Model CO-to-H2 conversion factors are derived for each set of physical conditions. We find the maximum value is 1.61020 cm-2/(K km s. Increasing [H2/12CO] to 105 to account for photodissociation and cosmic ray ionization increases the column density and upper limits by a factor of 10. Applying these limits to the CO luminosities, the upper limit on the total molecular mass in MW-C21 is 1322 M⊙, corresponding to 27% of the neutral gas mass. For the three clumps that are fully resolved, lower limits to the virial ratios are 28832, 6828, and 15739, which suggest that these structures are bound by external pressure to remain dynamically stable over the entrainment time of 2106 years or are being disrupted by shear and expansion over the clump crossing times of 3-8105 years. The observations presented in this study add to the growing census of cold gas entrained within the Galactic Center wind.
@@ 4.2 Clump masses
Using the RADEX models, an upper limit to the conversion factor, , can be derived that is consistent with the non-detection of 13CO emission. For each set of parameters, the model 12CO brightness temperature, TB(12CO), is calculated at the H2 column density upper limit. The model 12CO integrated intensity, , is the product of this brightness temperature and the line width of 11 km s-1. The upper limit on is =. The values for TB(12CO) and for each model are listed in Table 2.
@@ 4.3 Virial ratio
The molecular clumps are embedded within a cloud of atomic gas with a mass of 428 M⊙ distributed within a radius of 40 pc (Di Teodoro et al., 2018). Assuming a spherical cloud, the mean volume density is 0.06 cm-3. An upper limit on the gas kinetic temperature can be made assuming the measured HI 21 cm velocity dispersion from Di Teodoro et al. (2018) is due to purely thermal motions, such that =3.5104 K, where is the H i atomic mass and is the Boltzmann constant. The upper limit on the cloud thermal pressure, nT, is 2.2103 K cm-3= 4.410-13 dynes cm-2, which is insufficient to bound the internal pressure of the molecular clumps. Cosmic-ray pressure and magnetic field surface tension are possible bounding agents but the strength of each is not well defined in the local environment of MW-C21.
@@ 5 Conclusions
We have searched for CO emission from 2 atomic gas clouds associated with the Fermi Bubble. 12CO J=1-0 emission is detected in the cloud MW-C21. This emission is distributed in 11 discrete clumps. 12CO is not detected in the cloud MW-C20 and 13CO J=1-0 emission is not detected in either cloud. Based on RADEX models to predict 13CO J=1-0 brightness temperatures over a range of physical conditions, the absence of detected 13CO emission constrains the H2 column density to be less than 31021 cm-2 and the CO to H2 conversion factor, , to be 1.61020 cm-2(K km s-1)-1, assuming [H2/12CO]=104. Clump mass upper limits range from 1 M⊙ to 95 M⊙ with a total molecular mass 132 M⊙, which comprises 27% of the neutral gas mass in the cloud. Mass surface density upper limits for the three clumps that are well resolved by the 12″ beam are 18, 42, and 9 M⊙pc-2. The derived virial ratio lower limits for the three resolved clumps are high (288, 68, 157), yet molecular gas is still present over the entrainment time scale, which suggests that these molecular features are bound by an external pressure component.